OpenStudy (anonymous):

y=(x/sqrt.(a^2 - 1)) - (2/sqrt.(a^2 - 1)) * arctan(sinx/(a+ sqrt.(a^2 - 1) + cosx)))

OpenStudy (anonymous):

really, thats intresting

OpenStudy (anonymous):

prove that y' = 1/(a+cosx)

myininaya (myininaya):

\[y=(\frac{x}{\sqrt{a^2-1}}-\frac{2}{\sqrt{a^2-1}}) \cdot \tan^{-1}(\frac{\sin(x)}{a+\sqrt{a^2-1}}+\cos(x))\]?

OpenStudy (anonymous):

better format: \[y=\left(\begin{matrix}x \\ \sqrt{a ^{2}-1} \end{matrix}\right) - \left(\begin{matrix}2 \\ \sqrt{a^2 - 1}\end{matrix}\right)\arctan(\left(\begin{matrix}sinx \\ a + \sqrt{a^2+1} + cosx\end{matrix}\right)\]

myininaya (myininaya):

wait is my arctan part right are yours?

myininaya (myininaya):

or yours?*

myininaya (myininaya):

\[y=\frac{x}{\sqrt{a^2-1}}-\frac{2}{\sqrt{a^2-1}} \cdot \tan^{-1}(\frac{\sin(x)}{a+\sqrt{a^2-1}}+\cos(x))\]

OpenStudy (anonymous):

mine, the cosx should be under the sinx

OpenStudy (zarkon):

that's ugly ;)

myininaya (myininaya):

\[(\frac{x}{\sqrt{a^2-1}})'=\frac{1}{\sqrt{a^2-1}}(x)'=\frac{1}{\sqrt{a^2-1}}\] now we need to do the hard part find \[(-\frac{2}{\sqrt{a^2-1}} \tan^{-1}(\frac{\sin(x)}{a+\sqrt{a^2-1}+\cos(x)}))'\] \[-\frac{2}{\sqrt{a^2-1}} (\tan^{-1}(\frac{\sin(x)}{a+\sqrt{a^2-1}+\cos(x)}))'\] so to find the derivative of this term first I will say LET \[g=\tan^{-1}(\frac{\sin(x)}{a+\sqrt{a^2-1}+\cos(x)})\] now take tan( ) of both sides \[\tan(g)=\frac{\sin(x)}{a+\sqrt{a^2-1}+\cos(x)} (=\frac{opp}{adj})\]  so we need to find the hyp by using Pythagorean thm \[hyp=\sqrt{\sin^2(x)+(a+\sqrt{a^2-1}+\cos(x))^2}\] so anyways back to \[\tan(g)=\frac{\sin(x)}{a+\sqrt{a^2-1}+\cos(x)} \] we need to take derivative of both sides \[\sec^2(g) g'=\frac{\cos(x)(a+\sqrt{a^2-1}+\cos(x))-\sin(x)(0+0+\cos(x))} {(a+\sqrt{a^2-1}+\cos(x))^2}\] \[\sec^2(g) g'=\frac{\cos(x)[a+\sqrt{a^2-1}+\cos(x)-\sin(x)]}{(a+\sqrt{a^2-1}+\cos(x))^2}\] now to find g' we divide both sides by sec^2(g) \[g'=\frac{\cos(x)[a+\sqrt{a^2-1}+\cos(x)-\sin(x)]}{(\sec^2(g))(a+\sqrt{a^2-1}+\cos(x))^2}\] we can use our triangle to write sec^2(g) in terms of x \[g'=\frac{\cos(x)[a+\sqrt{a^2-1}+\cos(x)-\sin(x)]}{(\frac{(a+\sqrt{a^2-1}+\cos(x))}{\sqrt{\sin^2(x)+(a+\sqrt{a^2-1}+\cos(x))^2}})^2(a+\sqrt{a^2-1}+\cos(x))^2}\]